[figure: 9.7 cm]

Temperature dependence of saturation magnetostriction measured for F e_{81}Si_{3.5}B_{13.5}C_{2} amorphous films by a bending method based on

the Villari effect.

B. Kundys^{a}, Yu. Bukhantsev^{a}, H. Szymczak^{a}, M. R. J. Gibbs^{b} and R. Żuberek^{a} a Inst. of Physics of the Polish Academy of Sciences, Al. Lotników 32/46, 02-668, Warsaw, Poland

b Department of Physics and Astronomy, Univ. of Sheffield, Sheffield, S3 7RH, England

Magnetostriction measurements were performed at different temperatures for F e_{81}Si_{3.5}B_{13.5}C_{2} METGLAS\begin{matrix}r \\ ◯ \end{matrix} films using a strained substrate technique. The magnetostriction constant was deduced from magnetic properties measured using a vibrating sample magnetometer (VSM). The film strain was varied using a range of curved sample holders. From the change of magnetic anisotropy field as a function of controlled strain, the saturation magnetostriction constant can be determined with sufficient accuracy. The temperature dependence of magnetostriction of METGLAS\begin{matrix}r \\ ◯ \end{matrix} F e_{81}Si_{3.5}B_{13.5}C_{2} from 9 to 287K is reported. It is suggested that a single-ion mechanism of magnetostriction is dominant in this amorphous alloy.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : Bohdan Kundys

Address for correspondence : Institute of Physics Polish Academy of Sciences, Al. Lotników 32/46, 02-668, Warsaw, Poland

Email address : kundy@ifpan.edu.pl